### v2

# Failure Modes of Maximum Entropy RLHF

Abstract
In this paper, we show that Simple Preference Optimization (SimPO) can be derived as Maximum Entropy Reinforcement Learning, providing a theoretical foundation for this reference-free method. Motivated by SimPO’s strong performance in offline preference optimization, we investigate whether Maximum Entropy RL can achieve similar results in online RLHF settings. Our experiments find that Maximum Entropy RL consistently exhibits overoptimization and unstable KL dynamics, even at very low learning rates. Unlike KL-constrained methods that maintain stable training, entropy regularization fails to prevent reward hacking and appears to correlate with overoptimization. Lastly, we discuss possible explanations for why SimPO succeeds in offline settings while Maximum Entropy RL struggles in online scenarios. Our findings suggest that reference-free approaches may face distinct challenges when applied to online or offline preference learning.
Failure Modes of Maximum Entropy RLHF
Ömer Veysel Çağatan KUIS AI Center, Koç University Sariyer/˙Istanbul,Turkey ocagatan19@ku.edu.tr          Barış Akgün Koç Universityi KUIS AI Center Sariyer/˙Istanbul,Turkey baakgun@ku.edu.tr
1 Introduction
Aligning AI systems with human values is widely recognized as a central challenge in modern AI (bengio2025singaporeconsensusglobalai; stuart). The dominant paradigm, Reinforcement Learning from Human Feedback (RLHF) (christiano2023deepreinforcementlearninghuman; stiennon2022learningsummarizehumanfeedback; ziegler2020finetuninglanguagemodelshuman; bai2022traininghelpfulharmlessassistant; ouyang2022traininglanguagemodelsfollow), typically follows a three-stage pipeline: supervised fine-tuning, reward model training from preference data, and policy optimization with reinforcement learning under KL divergence regularization to constrain deviation from a reference model. While effective, this pipeline is computationally expensive and operationally complex, requiring separate reward models, substantial human annotation, and careful hyperparameter tuning to maintain stability.
These challenges have motivated direct alignment algorithms (DAAs) (rafailov2024scalinglawsrewardmodel), which bypass explicit reward modeling and online RL. Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) derives an analytical solution to a KL-regularized RL objective, expressing the reward implicitly as a function of the optimal policy and reducing preference learning to supervised optimization. More recently, Simple Preference Optimization (SimPO) (meng2024simposimplepreferenceoptimization) has demonstrated strong empirical performance while eliminating the reference model entirely, instead using length-normalized log-likelihoods and a fixed margin between preferred and dispreferred responses.
Despite its empirical success, SimPO has lacked a principled theoretical foundation comparable to that of reference-based methods such as DPO. In this work, we establish a connection between SimPO and Maximum Entropy Reinforcement Learning (ziebart2008maximum). We show that SimPO can be interpreted as a closed-form solution to a Maximum Entropy RL objective, providing a theoretical grounding analogous to DPO’s relationship with KL-constrained RL and suggesting that reference-free optimization can naturally arise from entropy regularization under appropriate conditions.
This perspective raises an important empirical question: if SimPO corresponds to an offline Maximum Entropy solution, can online Maximum Entropy RL serve as a viable alternative to KL-constrained methods in RLHF? To investigate this, we compare Maximum Entropy RL and KL-constrained RL on the TL;DR summarization benchmark (stiennon2022learningsummarizehumanfeedback) using models from the Pythia suite (biderman2023pythiasuiteanalyzinglarge).
Our results reveal a clear asymmetry. While SimPO performs reliably in offline preference optimization, online Maximum Entropy RL frequently exhibits instability and overoptimization, even at conservative learning rates. We observe that increases in entropy often correlate with these instabilities, indicating that entropy regularization alone does not reliably prevent reward hacking and may, in some cases, exacerbate it. We hypothesize that SimPO benefits from implicit stabilizing mechanisms—such as dataset constraints and target margins—that partially substitute for the regularization provided by a reference model, whereas these protections are absent in online Maximum Entropy RL.
Our contributions are threefold. First, we provide a theoretical interpretation of SimPO as a Maximum Entropy Reinforcement Learning solution, situating it within established RL frameworks. Second, we empirically demonstrate that directly applying Maximum Entropy RL in online RLHF settings can lead to instability and overoptimization, highlighting limitations of entropy regularization in isolation. Third, we offer insight into why SimPO succeeds offline despite these challenges, emphasizing the role of implicit regularization through data constraints and margin-based objectives. Together, these findings clarify the relationship between entropy-based methods and preference optimization, and point to the need for additional regularization mechanisms for robust reference-free alignment in online settings.
2 Background
In this section, we review the relevant background topics, while additional related work is provided in Appendix A.
2.1 Canonical RLHF
We reiterate the standart RLHF pipeline as outlined in (ziegler2020finetuninglanguagemodelshuman) and subsequent works (stiennon2022learningsummarizehumanfeedback; bai2022traininghelpfulharmlessassistant; ouyang2022traininglanguagemodelsfollow). It consists of three main stages: (1) Supervised Fine-Tuning (SFT), (2) Reward Modeling, and (3) RL Optimization.
SFT:
A pre-trained LM is fine-tuned on task-specific high-quality data via supervised learning to obtain the initial policy $\pi^{\text{SFT}}$.
Reward Modeling:
Prompts $x$ are sampled, and $\pi^{\text{SFT}}$ generates answer pairs $(y_{1},y_{2})$. Human annotators indicate preferences $y_{w}\succ y_{l}\mid x$, assumed to reflect a latent reward function $r(x,y)$. A common approach is to model preferences with the Bradley-Terry (BT) model (19ff28b9-64f9-3656-ba40-08326a05748e):
$p(y_{1}\succ y_{2}\mid x)=\frac{e^{r^{*}(x,y_{1})}}{e^{r^{*}(x,y_{1})}+e^{r^{*}(x,y_{2})}}$ (1)
Given a dataset $\mathcal{D}={x^{(i)},y_{w}^{(i)},y_{l}^{(i)}}$, we learn a reward model $r_{\phi}$ by minimizing the binary classification loss:
$\mathcal{L}R$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\quad\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$
where $\sigma$ is the sigmoid function. In practice, $r_{\phi}$ is initialized from $\pi^{\text{SFT}}$ with a linear head, and reward outputs are normalized for stability.
RL Fine-Tuning:
Finally, the policy $\pi_{\theta}$ is optimized using the learned reward, constrained by a KL term to stay close to the reference policy $\pi_{\text{ref}}=\pi^{\text{SFT}}$:
$\mathcal{L}_{R}=-\mathbb{E}{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right],$
This prevents overoptimization and distributional shift. In practice, this objective is optimized with PPO (schulman2017proximalpolicyoptimizationalgorithms), using a reward defined as $r(x,y)=r_{\phi}(x,y)-\beta(\log(\pi_{\theta}(y\mid x))-\log(\pi_{\text{ref}}(y\mid x)))$.
2.2 Direct Preference Optimization
Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) has become a popular method for preference-based tuning. Unlike traditional approaches that train a separate reward model, DPO defines the reward directly in terms of the optimized policy:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$
[…]
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
By modeling preferences directly through policy ratios, DPO removes the need for an explicit reward model while remaining grounded in a probabilistic preference framework.
2.3 Simple Preference Optimization
Simple Preference Optimization (SimPO) (meng2024simposimplepreferenceoptimization) is a reference-free method for preference-based fine-tuning that aligns the reward used in training with the likelihood used at inference. Unlike DPO, SimPO eliminates the need for a reference policy by defining the reward as the length-normalized log-likelihood of the model output:
$r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)=$
$\frac{\beta}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i})$
This formulation ensures that the reward ranking $r(x,y_{w})>r(x,y_{l})$ aligns with the generation-time likelihood ranking $p_{\theta}(y_{w}\mid x)>p_{\theta}(y_{l}\mid x)$, which is often violated in DPO. SimPO also introduces a target margin $\gamma>0$ into the Bradley-Terry model to encourage separation between preferred and dispreferred responses:
$p(y_{w}\succ y_{l}\mid x)=\sigma\left(r(x,y_{w})-r(x,y_{l})-\gamma\right)$
This leads to the SimPO training objective:
[…]
$\left.\left.\quad-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\gamma\right)\right]$
3 SimPO is the Maximum Entropy RL
SimPO is a widely used preference alignment method, appreciated for its strong empirical performance and simplicity due to its reference-free objective. However, it lacks a theoretical foundation, unlike reference-based approaches such as DPO, which is derived from a KL-constrained RL objective. Recent work (liu2024understandingreferencepoliciesdirect) made the important observation that posterior probability rewards correspond to Maximum Entropy RL in their analysis of reference policies. Building on this insight, we establish the connection between this Maximum Entropy RL formulation and SimPO, showing that SimPO can be understood as Maximum Entropy RL with the addition of length normalization and target margins from preference learning.
3.1 Maximum Entropy RL
Maximum Entropy Reinforcement Learning (MaxEnt RL) augments the standard RL objective with an entropy term, encouraging policies that align with the soft value function (ziebart2008maximum; toussaint2009robot; rawlik2013stochastic; fox2015taming; o2016combining; abdolmaleki2018maximum; sac; mazoure2020leveraging; han2021max; zhang2025maximumentropymisleadspolicy). It is deeply connected to probabilistic inference (toussaint2009robot; rawlik2013stochastic; levine2018reinforcement) and supported by both stochastic inference (ziebart2010modeling; eysenbach2021maximum) and game-theoretic foundations (grunwald2004game; ziebart2010maximum; han2021max; kim2023adaptive). MaxEnt is often favored for promoting exploration (sac; hazan2019provably), smoothing optimization (ahmed2019understanding), and enabling robust decision-making (eysenbach2021maximum).
The general form of the Maximum Entropy Reinforcement Learning (MaxEnt RL) objective can be written as
$\pi^{\star}=\operatorname*{arg\,max}_{\pi}\;\mathbb{E}_{\tau\sim p^{\pi}(\tau)}$
$\left[\sum_{t=1}^{T}r(\mathbf{s_{t}},\mathbf{a_{t}})+\alpha\,\mathcal{H}_{\pi}[\mathbf{a_{t}}\mid\mathbf{s_{t}}]\right]$
where $\tau=(\mathbf{s_{1}},\mathbf{a_{1}},\mathbf{s_{2}},\mathbf{a_{2}},\dots,\mathbf{s_{T}},\mathbf{a_{T}})$ is a trajectory sampled under policy $\pi$, and $p^{\pi}(\tau)=p_{1}(\mathbf{s_{1}})\prod_{t=1}^{T}\pi(\mathbf{a_{t}}\mid\mathbf{s_{t}})\,p(\mathbf{s_{t+1}}\mid\mathbf{s_{t}},\mathbf{a_{t}})$ denotes the trajectory distribution induced by $\pi$. The term $\mathcal{H}_{\pi}[\mathbf{a_{t}}\mid\mathbf{s_{t}}]=-\int\pi(\mathbf{a_{t}}\mid\mathbf{s_{t}})\log\pi(\mathbf{a_{t}}\mid\mathbf{s_{t}})\,d\mathbf{a_{t}}$ represents the conditional entropy of the policy at each time step, and the temperature coefficient $\alpha$ controls the trade-off between reward maximization and policy stochasticity.
3.2 SimPO from Maximum Entropy RL
RLHF is commonly modeled as a contextual bandit problem, though some approaches treat it as a token-level MDP (rafailov2024rqlanguagemodel; xie2024exploratorypreferenceoptimizationharnessing). In this work, we adopt the contextual bandit view (Elwood_2023), under which the maximum entropy formulation aligns with KL-constrained objectives. The resulting objective is given as follows.
$\max_{\pi}\mathbb{E}_{x\sim D,y\sim\pi}[r(x,y)]+\alpha D_{\mathcal{H}}[\pi(y|x)]$ (2)
It is straightforward to show that optimal policy of the equation 2 (proof in Appendix  C) is as follows:
$\pi_{r}(y|x)=\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right)$ (3)
Following the analytical approach used in DPO’s derivation, we can rearrange this optimal policy equation to express the reward function in terms of the policy:
$r(x,y)=\alpha\log\pi_{r}(y|x)+\alpha\log Z(x)$ (4)
Now, applying this reparameterization to the Bradley-Terry preference model. For the ground-truth reward $r^{*}$ and corresponding optimal policy $\pi^{*}$, the preference probability becomes:
$p^{*}(y_{1}\succ y_{2}|x)=\sigma(r^{*}(x,y_{1})-r^{*}(x,y_{2}))$ (5)
Substituting our reparameterization:
$p^{*}(y_{1}$ $\succ y_{2}|x)$
$=\sigma\big(\alpha\log\pi^{*}(y_{1}|x)+\alpha\log Z(x)$
$\qquad-\alpha\log\pi^{*}(y_{2}|x)-\alpha\log Z(x)\big)$
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right)$
Crucially, the partition function $Z(x)$ cancels out, eliminating the need to compute it explicitly. This establishes a direct connection between the Bradley–Terry preference model and the optimal policy induced by Maximum Entropy RL. To recover the SimPO training objective used in practice, we introduce two additional components that are motivated by empirical and structural considerations in preference learning rather than by the Maximum Entropy RL derivation itself.
First, we apply length normalization by scaling the temperature with the response length. This is necessary in SimPO because, unlike KL-constrained methods such as DPO where the reference policy induces implicit regularization through cancellation of log-probabilities, SimPO optimizes raw log-likelihoods and is therefore prone to severe length exploitation without explicit normalization. Second, inspired by ${\psi}PO$ (azar2023generaltheoreticalparadigmunderstand), we incorporate a target reward margin $\gamma>0$ as a fixed substitute for the adaptive margin implicitly provided by the reference policy in DPO. These additions shape the optimization landscape and improve empirical stability, but do not alter the underlying connection between SimPO and Maximum Entropy RL established above. This leads to the SimPO objective for a parametric policy $\pi_{\theta}$:
$\mathcal{L}_{\text{SimPO}}$ $(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg[\log\sigma\bigg($
$\quad\frac{\alpha}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)$
$\quad-\frac{\alpha}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\bigg)\bigg]$
This derivation reveals that SimPO is equivalent to Maximum Entropy RL under the contextual bandit formulation with addition of length normalization and target margin augmentation, making explicit the theoretical connection that underlies SimPO’s design.
Figure 1: RLHF reward and entropy bonus during training for Pythia 1B with different entropy coefficients at learning rates 1e-6 (left) and 1e-7 (right). Win rates are reported in the legend for each entropy bonus coefficient setting.
Theoretical Guarantees.
Following the same theoretical framework as DPO, SimPO inherits analogous guarantees regarding representational completeness, equivalence class preservation, and consistency under the Bradley-Terry preference model. The detailed proofs and formal statements of these properties are provided in Appendix C.
4 Maximum Entropy RLHF
Having established the theoretical connection between SimPO and Maximum Entropy RL, we now turn to the online RLHF setting. Our goal is to evaluate whether Maximum Entropy RL can perform comparably to its KL-constrained counterpart when applied directly to preference optimization.
4.1 Experimental Setup and Methodology
We conduct experiments using 1B, 2.8B, and 6.9B parameter models from the Pythia suite (biderman2023pythiasuiteanalyzinglarge), trained with RLOO (ahmadian2024basicsrevisitingreinforcestyle) on the TL;DR summarization dataset (stiennon2022learningsummarizehumanfeedback). All experiments follow the RLHF training recipe described by huang2024nimplementationdetailsrlhf and are implemented using the TRL library (vonwerra2022trl). Model alignment is evaluated using simulated preference win rates computed with GPT-4o-mini (openai2024gpt4ocard) as a proxy evaluator, measured against reference summaries for TL;DR using greedy decoding unless otherwise stated. Our experimental protocol closely follows the setup of rafailov2024scalinglawsrewardmodel, enabling direct comparison with prior work.
Our aim is to study overoptimization phenomena in online RLHF, which are difficult to quantify reliably in more complex training setups involving larger models, diverse tasks, or heterogeneous reward signals. The TL;DR benchmark with Pythia models provides a controlled setting in which overoptimization can be measured consistently through preference win rates against fixed reference summaries, allowing systematic comparison across optimization objectives and model scales.
We adopt RLOO as a critic-free alternative to the standard RLHF pipeline while optimizing equivalent reward objectives. In the KL-constrained formulation, the reward is defined as
$r(x,y)=r_{\phi}(x,y)$
$-\beta\Big(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\Big),$
whereas in the Maximum Entropy formulation, the reward takes the form
$r(x,y)=r_{\phi}(x,y)-\alpha\log\pi_{\theta}(y|x).$ (6)
When applying Maximum Entropy RL to sequence generation, the entropy term can be trivially increased by producing longer responses. Unlike KL-constrained objectives, which include an explicit reference policy that naturally counteracts such length effects, the reference-free formulation lacks an inherent mechanism to penalize verbosity. We therefore employ length normalization in the entropy term to ensure that the reward contribution is comparable across responses of different lengths and to prevent systematic length exploitation during optimization. This yields the reward
$r(x,y)=r_{\phi}(x,y)-\frac{\beta}{|y|}\log\pi_{\theta}(y|x).$ (7)
Figure 2: RLHF reward and entropy bonus during training for Pythia 2.8B with different entropy bonus coefficients at learning rates 1e-7 (left) and 5e-8 (right). Win rates are reported in the legend for each entropy bonus coefficient setting.
We also explored extending our evaluation to additional preference datasets such as Anthropic-HH (bai2022traininghelpfulharmlessassistant). In this setting, we attempted to reproduce reported baselines using both the official DPO implementation and the TRL library. Across implementations, we consistently observed that DPO failed to outperform the SFT baseline, with performance often degrading during training. This behavior exhibited substantial variability across runs and implementations, making it difficult to establish stable and reproducible reference baselines for systematic analysis. We therefore focus our empirical study on the TL;DR benchmark, which provides more consistent training dynamics suitable for controlled investigation.
4.2 Results and Analysis
Online Maximum Entropy RLHF across model scales.
We evaluate Maximum Entropy RL in an online RLHF setting using Pythia models at three scales: 1B, 2.8B, and 6.9B parameters. Learning rates are chosen based on prior work, with $1\times 10^{-6}$ corresponding to a regime where KL-constrained RLHF performs well, and smaller learning rates following common practice in reference-free methods. Entropy coefficients are selected via grid search.
Across all model scales, Maximum Entropy RL is highly sensitive to the learning rate. At $1\times 10^{-6}$, all models rapidly enter overoptimized regimes regardless of the entropy coefficient. Lower learning rates improve stability, but do not fully eliminate this behavior.
For Pythia 1B, reducing the learning rate to $1\times 10^{-7}$ yields stable training and improved win rates relative to the SFT baseline (Figure 1). However, these gains are not driven by entropy regularization: similar performance is achieved even when the entropy coefficient is set to zero. Moreover, stable runs exhibit decaying entropy bonuses, while overoptimized runs show increasing entropy bonuses, suggesting that entropy correlates with reward hacking rather than preventing it.
Scaling behavior.
For Pythia 2.8B, we restrict experiments to learning rates of $1\times 10^{-7}$ and $5\times 10^{-8}$, as higher learning rates consistently caused rapid overoptimization. Despite these conservative settings, all Maximum Entropy variants overoptimize and fail to outperform the SFT baseline (Figure 2). In contrast, KL-constrained RLHF remains stable and effective at a higher learning rate of $1\times 10^{-6}$.
Interestingly, Pythia 6.9B does not exhibit behavior intermediate between smaller and larger models; instead, its training dynamics more closely resemble those of Pythia 1B than Pythia 2.8B. Training remains stable at a learning rate of $1\times 10^{-7}$ (Figure 11), but increasing the learning rate to $1\times 10^{-6}$ (Figure 10) leads to overoptimization, mirroring the instability observed in smaller models. However, KL-constrained RLHF remains stable for Pythia 6.9B while consuming only a small KL budget (Figure 12), indicating that effective KL budgets vary across model scales and are not reliably controlled by reference-free objectives.
KL budget and optimization dynamics.
To better understand these failures, we analyze KL divergence during training. Well-tuned KL-constrained RLHF exhibits an initial growth phase followed by slow KL increase, keeping the policy close to the reference model. This behavior is sensitive to the KL coefficient, but provides a reliable mechanism for controlling optimization.
In contrast, Maximum Entropy RL exhibits fragile KL behavior across model scales. For Pythia 1B and 6.9B, KL divergence grows approximately linearly without immediate collapse, whereas for Pythia 2.8B similar KL magnitudes correspond to severe overoptimization. This indicates that effective KL budgets are inherently model-dependent, and that reference-free objectives lack a mechanism to infer or enforce an appropriate optimization budget. As a result, their success is not predictable across model scales.
Figure 3: KL divergence metrics for KL-constrained and Maximum Entropy RL across training. Top row shows KL divergence between the current policy and the reference policy ($KL(\pi_{t}\|\pi_{\text{ref}})$) for Pythia 1B (left) and 2.8B (right). Bottom row shows KL divergence between consecutive policy iterations ($KL(\pi_{t}\|\pi_{t-1})$).
KL update magnitudes and optimization stability.
To disentangle whether overoptimization arises from the objective or from policy optimization dynamics, we examine the KL divergence between consecutive policy updates. Since RLOO implements policy optimization using PPO, one might attribute instability in Maximum Entropy RL to PPO-specific issues, such as poor ratio control or clipping behavior.
However, Figures  3,  10,  11 show that this explanation is insufficient. In KL-constrained runs, PPO reliably maintains bounded KL updates between successive policies, even in regimes that eventually overoptimize when the KL coefficient is relaxed. In contrast, Maximum Entropy runs consistently exhibit increasing KL drift between updates, including in runs that appear well-behaved early in training. This effect intensifies over time despite using substantially lower learning rates than those required for stable KL-constrained optimization.
These observations suggest that the instability is not merely algorithmic, but objective-driven. Without an explicit reference policy anchoring the optimization, reward gradients become sharper and induce more aggressive parameter updates, leading to compounding KL drift. Tightening PPO clipping alone is insufficient to prevent this behavior: for Pythia 2.8B, we experimented with clipping ranges as small as $10^{-4}$, yet still observed increasing KL drift and eventual overoptimization. This indicates that reference-free objectives lack the structural safeguards needed to control update magnitudes in online RLHF.
4.2.1 Minimum Entropy RL
Motivated by the link between maximum entropy and overoptimization and recent work showing entropy minimization can serve as an effective reward signal for LLM reasoning (agarwal2025unreasonableeffectivenessentropyminimization), we adopt an unconventional strategy: minimizing entropy to discourage excessively high-entropy which we expect to prevent overoptimization.
Our experiments reveal that Minimum Entropy RL prevents overoptimization and achieves competitive performance with Pythia-1B even at the same learning rate used by KL-constrained RL, under which Maximum Entropy collapses. Yet, with Pythia-2.8B, entropy minimization proves unstable: it is either too conservative, stalling learning, or too loose, leading to overoptimization. While entropy minimization succeeds as a standalone reward for reasoning, combining it with preference-based rewards appears to create optimization instabilities. Reducing the learning rate might offer some improvement, but Minimum Entropy is not a one-to-one substitute for KL, which remains more dynamic and adaptive. Lastly, this underscores that reference-free methods break down once they move outside a healthy KL budget, limiting their reliability.
4.3 Offline Maximum Entropy RLHF (SimPO)
Although Maximum Entropy RL fails to provide sufficient regularization in online settings, its closed-form offline solution, SimPO, is both effective and empirically competitive. This effectiveness cannot be attributed solely to conservative optimization: for example, one SimPO configuration for Llama 3 (grattafiori2024llama3herdmodels) uses a higher learning rate than DPO. Nevertheless, maintaining a sufficiently low learning rate remains critical for controlling the KL divergence, which appears to be a universal requirement across alignment methods (gao2022scalinglawsrewardmodel; rafailov2024scalinglawsrewardmodel).
A common argument for offline stability is that all training samples remain in-distribution, removing the need for explicit out-of-distribution regularization. However, as noted by azar2023generaltheoreticalparadigmunderstand; rafailov2024scalinglawsrewardmodel, the reward model can effectively drift out of distribution during optimization, leading to sub-epoch overoptimization. This suggests that stability depends not only on data coverage, but also on the structure of the reward objective itself. To address this, huang2025correctingmythosklregularizationdirect propose replacing KL regularization with $\chi^{2}$ regularization to inject pessimism directly into the reward model, reporting improved stability across multiple epochs. We were unable to reproduce these results, leaving open the question of whether $\chi^{2}$ regularization reliably implements pessimism. In this context, SimPO’s performance is particularly intriguing, not because it enforces pessimism, but because it achieves strong results without relying on a reference model.
To better understand this behavior, consider the pairwise reward used in DPO:
$r(y_{w}|x)-r(y_{l}|x)$ $=\beta\Bigg(\log\frac{\pi(y_{w}|x)}{\pi(y_{l}|x)}$
$\quad-\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\Bigg),$
where the second term reflects the contribution of the reference model. Because both $y_{w}$ and $y_{l}$ are sampled from the reference distribution, this term is typically negative but small, acting as an adaptive regularizer. This behavior parallels the role of the target margin in SimPO, with the key distinction that SimPO employs a fixed margin rather than a reference-dependent one (ahrabian2025practicalanalysishumanalignment).
Figure 4: Batch average of $\log\!\left(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ during DPO training on Pythia 1B.
To examine this connection empirically, we visualize the reference log-probability margins
$\log\!\left(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
during DPO training with Pythia 1B in Figure 4. We observe that these margins lie within a relatively narrow range, consistent with the fixed margins commonly used in reference-free methods such as SimPO. This suggests that, in offline settings, reference models may primarily serve to provide a bounded margin signal rather than strong distributional control.
Several caveats are worth noting. Large margins combined with high learning rates can drive aggressive separation, exacerbating reward overoptimization (rafailov2024scalinglawsrewardmodel). We observe extreme likelihood decreases indicative of such behavior, consistent with prior analyses of direct alignment algorithms. While reference models may implicitly adapt margins across samples—providing smaller corrections for easy examples and larger ones for harder cases—fixed margins lack this adaptivity and can amplify pathological updates under aggressive optimization. These findings suggest that reference models are neither necessary nor sufficient to prevent overoptimization in offline preference optimization. Instead, stability emerges from the interaction between learning rate, margin choice, and model capacity. We analyze these dynamics in more detail in Appendix B.
5 Conclusion
This work establishes a theoretical foundation for SimPO by connecting it to Maximum Entropy Reinforcement Learning, while revealing a striking asymmetry between offline and online performance. Although SimPO excels in offline preference optimization, our empirical investigation shows that online Maximum Entropy RL suffers from instability and overoptimization, with entropy regularization paradoxically correlating with rather than preventing reward hacking. These findings highlight that reference-free approaches, while appealing for their simplicity, may face fundamental limitations in online training scenarios, and suggest that SimPO’s success stems from implicit stabilizing factors such as dataset constraints and target margins that approximate the regularization benefits of reference models.
Acknowledgements
This work was funded by KUIS AI Center, Koç University.
Appendix A Additional Related Work
Reference-free Alignment.
[…]
Overoptimization in Preference Learning
Reward hacking (skalse2025definingcharacterizingrewardhacking) is a long-standing problem in reinforcement learning (Sutton1998) where policies achieve high rewards but fail to meet the actual objective (amodei2016concreteproblemsaisafety; hadfieldmenell2020inverserewarddesign; pan2022effectsrewardmisspecificationmapping). In language model alignment, this manifests as models learning to generate outputs that score highly on proxy metrics while being of poor actual quality. This overoptimization phenomenon was first systematically studied in traditional RLHF (christiano2023deepreinforcementlearninghuman; stiennon2022learningsummarizehumanfeedback; gao2022scalinglawsrewardmodel; ouyang2022traininglanguagemodelsfollow), where optimizing imperfect proxy reward models leads to qualitatively worse outputs, including overly wordy responses and hallucinated information.
[…]
Appendix B Margins and Overoptimization
It has been shown that methods such as SimPO can achieve performance comparable to DPO even with a target margin of $\gamma=0$, as demonstrated in the original SimPO paper. This suggests that offline methods do not necessarily require reference models when operating within the safe KL region, and that introducing margins generally improves performance across benchmarks. This effect arises from both model capabilities and dataset coverage: larger models are less prone to common overfitting behaviors and can extract more meaningful signals during optimization, rather than engaging in reward hacking, a phenomenon observed in both online and offline preference optimization (gao2022scalinglawsrewardmodel; rafailov2024directpreferenceoptimizationlanguage). Consequently, the influence of the reference model is minimal and can often be neglected. However, this behavior is contingent on the task being sufficiently challenging and the model being strong enough to avoid overoptimization. To validate this observation, we train Pythia-1B on TL;DR using SimPO across different margin values ($\gamma$) and learning rates, in a setting where the model is relatively weaker and the task is easier compared to standard chat datasets such as UltraFeedback (cui2024ultrafeedbackboostinglanguagemodels) used in SimPO.
We first consider a learning rate of $1\times 10^{-6}$, which is known to be effective for DPO, DPO metrics in Figure 9. In this setting, all SimPO models exhibit overoptimization regardless of the $\gamma$ hyperparameter, SimPO metrics in Figure  8 . Although reward definitions differ and direct comparison of losses or other training metrics is challenging, log-probabilities of samples remain comparable. We observe the characteristic extreme likelihood decreases, which correlate with overoptimization; this pattern is present in DAAs and, as we show, also occurs in online methods. Increasing the margin exacerbates this issue, as optimization aggressively seeks high separation, naturally resulting in overoptimization.
Reference-free methods like SimPO are particularly susceptible because they lack prior knowledge about sample difficulty, treating all samples equally. Some samples are inherently harder and should receive more attention, a behavior that could be partially captured by negative reference contributions in pairwise preferences. When a hardcoded margin pushes the model to satisfy strict separation objectives, it can amplify pathological behaviors during training.
However, when using a relatively low learning rate that allows for gradual updates, SimPO performs significantly better, metrics in Figure  7 and win rates in Figure  6 . In this regime, it emerges as a strong preference optimization method: an appropriate margin encourages the model to learn and optimize meaningful signals. Therefore, reference-free models require extra safeguards against overoptimization. Controlling the learning rate can act as an anchor, keeping updates within meaningful distributional shifts, although these models can still experience the overoptimization patterns observed in DAAs.
Appendix C Mathematical Derivations for Maximum Entropy RL
C.1 Deriving the Optimum of the Entropy-Regularized Reward Maximization Objective
In this appendix, we will derive the optimal policy for Maximum Entropy RL. Analogously to the KL-constrained case (rafailov2024directpreferenceoptimizationlanguage), we optimize the following objective:
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi}\bigl[r(x,y)\bigr]+\alpha\mathcal{H}\bigl[\pi(y|x)\bigr]$
under any reward function $r(x,y)$ and a general non-parametric policy class, where $\mathcal{H}[\pi(y|x)]=-\mathbb{E}_{y\sim\pi(y|x)}[\log\pi(y|x)]$ is the entropy of the policy. We now have:
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi}\bigl[r(x,y)\bigr]+\alpha\mathcal{H}\bigl[\pi(y|x)\bigr]$
$=\max_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y|x)}\left[r(x,y)-\alpha\log\pi(y|x)\right]$
$=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y|x)}\left[\log\pi(y|x)-\frac{1}{\alpha}r(x,y)\right]$
$=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y|x)}\Bigg[\log\frac{\pi(y|x)}{\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right)}$
$\qquad\qquad\qquad\qquad-\log Z(x)\Bigg]$
where we have partition function:
$Z(x)=\sum_{y}\exp\left(\frac{1}{\alpha}r(x,y)\right).$
Note that the partition function is a function of only $x$ and the reward function $r$, but does not depend on the policy $\pi$. We can now define
$\pi^{*}(y|x)=\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right),$
which is a valid probability distribution as $\pi^{*}(y|x)\geq 0$ for all $y$ and $\sum_{y}\pi^{*}(y|x)=1$. Since $Z(x)$ is not a function of $y$, we can then re-organize the final objective in Eq. C.1 as:
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\bigg[\mathbb{E}_{y\sim\pi(y|x)}\left[\log\frac{\pi(y|x)}{\pi^{*}(y|x)}\right]$
$\qquad\qquad\qquad-\log Z(x)\bigg]=$
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[D_{\text{KL}}(\pi(y|x)\|\pi^{*}(y|x))-\log Z(x)\right]$
Since $Z(x)$ is independent of $\pi$, the minimum is attained by the policy that minimizes the first KL term. By Gibbs’ inequality, the KL divergence reaches its minimum value of zero if and only if the two distributions are identical. Therefore, this yields the optimal solution:
$\pi(y|x)=\pi^{*}(y|x)=\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right)$ (8)
for all $x\in\mathcal{D}$. This completes the derivation.
C.2 Deriving the SimPO Objective Under the Bradley-Terry Model
It is straightforward to derive the SimPO objective under the Bradley-Terry preference model as we have
$p^{*}(y_{1}\succ y_{2}|x)=\frac{\exp\left(r^{*}(x,y_{1})\right)}{\exp\left(r^{*}(x,y_{1})\right)+\exp\left(r^{*}(x,y_{2})\right)}$ (9)
We can express the (unavailable) ground-truth reward through its corresponding optimal policy:
$r^{*}(x,y)=\alpha\log\pi^{*}(y|x)+\alpha\log Z(x)$ (10)
Substituting Eq. C.2 into Eq. C.2 we obtain:
$p^{*}(y_{1}$ $\succ y_{2}|x)$
$=\frac{\exp(\alpha\log\pi^{*}(y_{1}|x))}{\sum_{i\in\{1,2\}}\exp(\alpha\log\pi^{*}(y_{i}|x))}$
$=\frac{1}{1+\exp\left(\alpha\log\frac{\pi^{*}(y_{2}|x)}{\pi^{*}(y_{1}|x)}\right)}$
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right)$
The last line is the per-instance loss for SimPO, without target margin $\gamma$ and length normalization.
C.3 Deriving the SimPO Objective Under the Plackett-Luce Model
The Plackett-Luce model (af5079a1-8ca5-3727-a405-0a82390327b7) extends the Bradley-Terry model from pairwise comparisons to full rankings. As in the Bradley-Terry framework, the probability of selecting an option is assumed to be proportional to the value of an underlying latent reward function. In our setting, given a prompt $x$ and a collection of $K$ candidate answers $y_{1},\ldots,y_{K}$, the user produces a permutation $\tau:[K]\to[K]$ that represents their ranking of the answers. Under the Plackett-Luce model, the probability of such a ranking is defined as follows. Let $r_{k}=r^{*}(x,y_{\tau(k)})$. Then:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp(r_{k})}{\sum_{j=k}^{K}\exp(r_{j})}$ (11)
Observe that when $K=2$, Equation 11 simplifies to the Bradley-Terry model. For the general Plackett-Luce model, however, we can still leverage the reward parameterization by substituting the reward function expressed in terms of its optimal policy. As in Appendix C.2, the normalization constant $Z(x)$ cancels out. Let $s_{k}=\alpha\log\pi^{*}(y_{\tau(k)}|x)$. Then:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp(s_{k})}{\sum_{j=k}^{K}\exp(s_{j})}$ (12)
Similarly to the approach for standard DPO, if we have access to a dataset $\mathcal{D}=\{\tau^{(i)},y_{1}^{(i)},\ldots,y_{K}^{(i)},x^{(i)}\}_{i=1}^{N}$ of prompts and user-specified rankings, we can use a parameterized model and optimize this objective with maximum-likelihood. Let $s_{k}^{\theta}=\alpha\log\pi_{\theta}(y_{\tau(k)}|x)$. Then:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})$
$=-\mathbb{E}_{\mathcal{D}}\left[\sum_{k=1}^{K}\log\frac{\exp(s_{k}^{\theta})}{\sum_{j=k}^{K}\exp(s_{j}^{\theta})}\right]$
C.4 Deriving the Gradient of the SimPO Objective
In this section we derive the gradient of the SimPO objective:
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})$
$=-\nabla_{\theta}\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\alpha\log\pi_{\theta}(y_{w}|x)-\alpha\log\pi_{\theta}(y_{l}|x)\right)\right]$
We can rewrite the RHS of Equation C.4 as
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\frac{\sigma^{\prime}\left(u\right)}{\sigma\left(u\right)}\nabla_{\theta}\left(u\right)\right],$
where $u=\alpha\log\pi_{\theta}(y_{w}|x)-\alpha\log\pi_{\theta}(y_{l}|x)$.
Using the properties of sigmoid function $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$ and $\sigma(-x)=1-\sigma(x)$, we obtain the final gradient
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\\ \bigg[\alpha\sigma\left(\alpha\log\pi_{\theta}(y_{l}|x)-\alpha\log\pi_{\theta}(y_{w}|x)\right)\\ \bigg[\nabla_{\theta}\log\pi(y_{w}\mid x)-\nabla_{\theta}\log\pi(y_{l}\mid x)\bigg]\bigg],$
After using the reward substitution of $\hat{r}_{\theta}(x,y)=\alpha\log\pi_{\theta}(y\mid x)$ we obtain the final form of the gradient.
C.5 Proof of Lemma 1 and 2 from DPO for Maximum Entropy RL
In this section, we will prove the two lemmas from DPO for Maximum Entropy RL.
Lemma 1 (Lemma 1).
Under the Plackett-Luce preference framework, and in particular the Bradley-Terry framework, two reward functions from the same equivalence class induce the same preference distribution.
Proof.
We say that two reward functions $r(x,y)$ and $r^{\prime}(x,y)$ are from the same equivalence class if $r^{\prime}(x,y)=r(x,y)+f(x)$ for some function $f$. We consider the general Plackett-Luce (with the Bradley-Terry model a special case for $K=2$) and denote the probability distribution over rankings induced by a particular reward function $r(x,y)$ as $p_{r}$. For any prompt $x$, answers $y_{1},\ldots,y_{K}$ and ranking $\tau$ we have: Let $r_{k}=r(x,y_{\tau(k)})$ and $r^{\prime}_{k}=r^{\prime}(x,y_{\tau(k)})=r_{k}+f(x)$. Then: Let $r_{k}=r(x,y_{\tau(k)})$ and $r^{\prime}_{k}=r_{k}+f(x)$. Then:
$p_{r^{\prime}}$ $(\tau|y_{1},\ldots,y_{K},x)$
$=\prod_{k=1}^{K}\frac{\exp(r^{\prime}_{k})}{\sum_{j=k}^{K}\exp(r^{\prime}_{j})}$
$=\prod_{k=1}^{K}\frac{\exp(r_{k}+f(x))}{\sum_{j=k}^{K}\exp(r_{j}+f(x))}$
$=\prod_{k=1}^{K}\frac{\exp(f(x))\exp(r_{k})}{\exp(f(x))\sum_{j=k}^{K}\exp(r_{j})}$
$=\prod_{k=1}^{K}\frac{\exp(r_{k})}{\sum_{j=k}^{K}\exp(r_{j})}$
$=p_{r}(\tau|y_{1},\ldots,y_{K},x)$
which completes the proof. ∎
Lemma 2 (Lemma 2).
Two reward functions from the same equivalence class induce the same optimal policy under the entropy-regularized RL problem.
Proof.
Let us consider two reward functions from the same class, such that $r^{\prime}(x,y)=r(x,y)+f(x)$ and, let us denote as $\pi_{r}$ and $\pi_{r^{\prime}}$ the corresponding optimal policies. For all $x,y$, let $\rho=\frac{1}{\alpha}r(x,y)$, $\rho^{\prime}=\frac{1}{\alpha}r^{\prime}(x,y)$, and $\phi=\frac{1}{\alpha}f(x)$. Then:
$\pi_{r^{\prime}}(y|x)$ $=\frac{\exp(\rho^{\prime})}{\sum_{y}\exp(\rho^{\prime})}$
$=\frac{\exp(\rho+\phi)}{\sum_{y}\exp(\rho+\phi)}$
$=\frac{\exp(\phi)\exp(\rho)}{\exp(\phi)\sum_{y}\exp(\rho)}$
$=\frac{\exp(\rho)}{\sum_{y}\exp(\rho)}$
$=\pi_{r}(y|x),$
C.6 Proof of Theorem 1 from DPO for Maximum Entropy RL
In this section, we will eloborate on the results of the main theorem from DPO for Maximum Entropy RL.
Theorem 1 (Maximum Entropy Version).
Assume we have a parameter $\alpha>0$. All reward equivalence classes, as defined in the previous section, can be represented with the reparameterization $r(x,y)=\alpha\log\pi(y|x)$ for some model $\pi(y|x)$.
Proof.
Consider any reward function $r(x,y)$, which induces an optimal model $\pi_{r}(y|x)$ under the entropy-regularized RL problem, with solution given by the optimal policy derivation. We have:
$r(x,y)=\alpha\log\pi_{r}(y|x)+\alpha\log Z(x)$
where $Z(x)=\sum_{y}\exp\left(\frac{1}{\alpha}r(x,y)\right)$ (notice that $Z(x)$ also depends on the reward function $r$). Using the operator $r^{\prime}(x,y)=f(r,\alpha)(x,y)=r(x,y)-\alpha\log Z(x)$, we see that this new reward function is within the equivalence class of $r$ and, we have:
$r^{\prime}(x,y)=\alpha\log\pi_{r}(y|x)$
which completes the proof. ∎
We can further expand on these results. We can see that if $r$ and $r^{\prime}$ are two reward functions in the same class, then
$f(r,\alpha)(x,y)=\alpha\log\pi_{r}(y|x)=$
$\alpha\log\pi_{r^{\prime}}(y|x)=f(r^{\prime},\alpha)(x,y)$
where the second equality follows from Lemma 2. We have proven that the operator $f$ maps all reward functions from a particular equivalence class to the same reward function. Next, we show that for every equivalence class of reward functions, the reward function that has the reparameterization outlined in the main theorem is unique.
Proposition 1.
Assume we have a parameter $\alpha>0$. Then every equivalence class of reward functions has a unique reward function $r(x,y)$, which can be reparameterized as $r(x,y)=\alpha\log\pi(y|x)$ for some model $\pi(y|x)$.
Proof.
We will proceed using proof by contradiction. Assume we have two reward functions from the same class, such that $r^{\prime}(x,y)=r(x,y)+f(x)$. Moreover, assume that $r^{\prime}(x,y)=\alpha\log\pi^{\prime}(y|x)$ for some model $\pi^{\prime}(y|x)$ and $r(x,y)=\alpha\log\pi(y|x)$ for some model $\pi(y|x)$, such that $\pi\neq\pi^{\prime}$. We then have
$r^{\prime}(x,y)=r(x,y)+f(x)=\alpha\log\pi(y|x)+f(x)$
$=\alpha\log\pi(y|x)\exp(\frac{1}{\alpha}f(x))$
$=\alpha\log\pi^{\prime}(y|x)$
for all prompts $x$ and completions $y$. Then we must have $\pi(y|x)\exp(\frac{1}{\alpha}f(x))=\pi^{\prime}(y|x)$. Since these are distributions, summing over $y$ on both sides, we obtain that $\exp(\frac{1}{\alpha}f(x))=1$ and since $\alpha>0$, we must have $f(x)=0$ for all $x$. Therefore $r(x,y)=r^{\prime}(x,y)$. This completes the proof. ∎
We have now shown that every reward class has a unique reward function that can be represented as outlined in the main theorem, which is given by $f(r,\alpha)$ for any reward function in that class.
Appendix D Extra Figures
Figure 5: KL divergence evolution during training for 1B and 2.8B parameter models using different regularization methods. The left panel shows results for the 1B model and the right panel shows results for the 2.8B model. Each panel compares KL-Constrained and Maximum-Entropy approaches. Checkmarks ($\checkmark$) indicate high win rate runs and crosses ($\times$) indicate overoptimized runs.
Figure 6: Win rate progression across training checkpoints for different values of the gamma hyperparameter. Results are for the Pythia-1B model trained with a learning rate of $2\times 10^{-7}$.
Figure 7: SimPO training metrics across different gamma values. Comparison of key training dynamics including loss, gradients, logits, and reward metrics for $\gamma\in\{0,0.25,0.5,1.0\}$ using Pythia-1B with learning rate $2\times 10^{-7}$.
Figure 8: SimPO training metrics across different gamma values. Comparison of key training dynamics including loss, gradients, logits, and reward metrics for $\gamma\in\{0,0.25,0.5,1.0,2.0\}$ using Pythia-1B with learning rate $1\times 10^{-6}$.
Figure 9: DPO training metrics with $\beta=0.05$. Comparison of key training dynamics including loss, gradients, logits, and reward metrics, using Pythia-1B with learning rate $1\times 10^{-6}$.
Figure 10: Reward dynamics and KL divergence metrics for entropy-regularized RL training across different entropy coefficients. Top-left panel shows reward progression (RLHF reward) over training steps for various entropy values. Top-right panel shows KL divergence between the current policy and the SFT reference policy ($KL(\pi_{t}||\pi_{\text{SFT}})$). Bottom-left panel tracks entropy reward across training steps. Bottom-right panel displays KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All plots are based on the Pythia-6.9B model trained with the learning rate of $1\times 10^{-6}$.
Figure 11: Reward dynamics and KL divergence metrics for entropy-regularized RL training across different entropy coefficients. Top-left panel shows reward progression (RLHF reward) over training steps for various entropy values. Top-right panel shows KL divergence between the current policy and the SFT reference policy ($KL(\pi_{t}||\pi_{\text{SFT}})$). Bottom-left panel tracks entropy reward across training steps. Bottom-right panel displays KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All plots are based on the Pythia-6.9B model trained with the learning rate of $1\times 10^{-7}$.
Figure 12: Training metrics for KL-constrained RL on the Pythia-6.9B model. Top panel shows the KL divergence between the policy and reference SFT policy ($KL(\pi_{t}||\pi_{\text{SFT}})$) over training steps. Middle panel displays the reward trajectory (RLHF reward). Bottom panel shows the KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All results correspond to a single training run with a fixed KL constraint.